a movie theater $7.50 for each adult ticket and $4 for each child ticket. if the theater sold a total of 145 tickets for a total of $790, how many adult tickets were sold
Answer by mananth(16949)

x= number of adult tickets
y= number of child tickets

1.00 x + 1.00 y = 145.00 .............1
Total value
7.50 x + 4.00 y = 790.00 .............2
Eliminate y
multiply (1)by -4.00
Multiply (2) by 1.00
-4.00 x -4.00 y = -580.00
7.50 x + 4.00 y = 790.00
Add the two equations
3.50 x = 210.00
/ 3.50
x = 60.00
plug value of x in (1)
1.00 x + 1.00 y = 145.00
60.00 + y = 145.00
y = 145.00 -60.00
y = 85.00
y = 85.00
x= 60 number of adult tickets
y= 85 number of child tickets
m.ananth@hotmail.ca